OpenStudy (aravindg):

A 100km/hr wind blows normally against one wall of a house with an area of 50 m/s^2 .Calculate the force exerted on the wall if the air moves parallel to the wall after striking it and has a density of 1.134 kg/m^3.

OpenStudy (aravindg):

@Expert , @Vincent-Lyon.Fr

OpenStudy (aravindg):

@experimentX

OpenStudy (vincent-lyon.fr):

F = 43750 N ?

OpenStudy (aravindg):

ys @Vincent-Lyon.Fr hw did u get it?

OpenStudy (vincent-lyon.fr):

\(F=\rho Av^2\)

OpenStudy (aravindg):

just put in values?

OpenStudy (vincent-lyon.fr):

Yes, with v in m/s, not in km/h and A in m²

OpenStudy (mos1635):

there is no need to derive taht formula??

OpenStudy (aravindg):

if u can do it @mos1635

OpenStudy (mos1635):

do you need it? it's not simple....

OpenStudy (aravindg):

tell anyway

OpenStudy (mos1635):

ok are you familiar with ΣF=ΔP/Δt ??

OpenStudy (aravindg):

ys

OpenStudy (aravindg):

newton 2nd law

OpenStudy (mos1635):

ok each molucule of air bounce back from wall (assume elastic collition) the for 1 molucule ΔP=-2*m*v so for wall ΔP=2*m*v at time Δt there will be N molucules that hit wall  those N molucules will be in volume ΔV=A*d d=v*Δt and if N'= density in molucules /m^3 then N=N'*ΔV N=N'*A*v*Δt so F(total)=ΔP(total)/Δt Ft=N*(2*m*v)/Δt Ft=N'*A*v*Δt*(2*m*v)/Δt Ft=N'*A*v*(2*m*v) Ft=2*N'*m*A*v^2 Ft=2 * ρ * A * v^2 (that factor of 2 is messing thing up :) :) )